What is the morality of a nitric acid solution when 60.5 mL of HNO3 are required to react with 25.0 mL of a 1.00 M Barium Hydroxide Solution? (Hint: A Balanced Reaction is needed)
1
Expert's answer
2020-04-02T08:56:44-0400

Ba(OH)2 + 2HNO3 = Ba(NO3)2 + 2H2O

C(Ba(OH)2)*V(Ba(OH)2) = C(HNO3)*V(HNO3)

C(HNO3) = C(Ba(OH)2)*V(Ba(OH)2)/V(HNO3) = 1M*25mL/60.5mL = 0.413 M